# Fourier PINNs: From Strong Boundary Conditions to Adaptive Fourier Bases

### v1

@@ 6 Experiments
In this section, we present a series of tests comparing our methods against a range of PDE problems. These include the 1D and 2D Poisson equations, which are canonical linear elliptic PDE problems, and the 1D and 2D (steady-state) Allen-Cahn equations, which are non-linear elliptic PDEs demonstrating the effect of adding a non-linear reaction term. Additionally, we analyze the 1D (one-way) Wave equation to observe the impact of incorporating a first time derivative. We first describe the baseline methods we compare against Fourier PINNs, then list our experimental details. Finally, we provide a detailed discussion after presenting the results comparing each baseline on the example problems.
@@ 6.1 Baselines
We benchmarked against the following state-of-the-art PINN models for solving high-frequency and/or multi-scale solutions:
@@
    (RFF-PINN) Random Fourier Feature PINNs (Wang et al., 2021) with dynamically re-weighted loss terms derived from the NTK eigenvalues as in Wang et al. (2022b),
@@
    (W-PINN) Weighted PINNs that down-weight the residual loss term to reduce its dominance and to better fit the boundary loss Wang et al. (2022b);
@@
    (Spectral) standard Spectral Methods (Boyd, 2001), which approximate solutions using a linear combination of Fourier bases, and estimates the basis coefficients via least mean squares [^2] Throughout the experiments, we used the same set of Fourier bases for the spectral method and our method..
We do not test against the conceptually similar Fourier Neural Operators Li et al. (2020) as they are designed to solve the inverse problem, thus requiring a data loss term. Our method is designed to solve forward problems and requires no training data within the domain.
Table 1: This table lists the number and corresponding scales of Gaussian variances utilized in the RFF-PINN. Scales were selected randomly from a set range of $[1,200]$. Preference was given to the subset $[1,20,50,100]$ to ensure a balanced distribution across the available range.
@@ 6.2 Hyperparameters and experimental details
All code is implemented in PyTorch [^3] Code will be made available upon acceptance. For all the NN based methods, we used two hidden layers, with $100$ neurons per layer and the Tanh activation function. We ran every method for five random trials, and report the five number summaries for each test in boxplot form. For PINN, W-PINN, A-PINN and RFF-PINN, we first ran $40$K iterations of the Adam optimizer (Kingma & Ba, 2014) with a learning rate of $10^{-3}$, and then L-BFGS optimizer (Liu & Nocedal, 1989) until convergence (with the tolerance set to $10^{-9}$). In Fourier PINNs, we specified the range of frequency candidates in the Fourier bases by setting a maximum frequency $K$ and including all equally spaced frequencies in the set $\{1,2,...,K\}$. In Algorithm 1, we set $\alpha=0.1$, the number of iterations $T$ to 1K, the inner iteration number $A$ to 5, the truncation threshold $\delta$ to $10^{-4}$, and the outer iteration number $E$ to 40. RFF-PINNs require specifying both the number and scales of the Gaussian variances used to construct the random features. To comprehensively evaluate the effect of both hyperparameters, we tested RFF-PINNs with 20 different settings. We used one, two, three and five scales, and set the variances to common values suggested in (Wang et al., 2021), combined with randomly sampled ones. The exact sets of chosen scales are listed in Table 1. For W-PINN, we varied the weight of the residual loss from $\{10^{-1},10^{-3},10^{-4}\}$, and for A-PINN, we introduced a learned parameter for the activation function in each layer and update these parameters jointly.
Figure 4: 1D Poisson with $u(x)=\sin(100x)$
Figure 5: 1D Poisson with $u(x)=\sin(x)+0.1\sin(20x)+0.05\cos(100x)$
Figure 6: 1D Poisson with $u(x)=\sin(6x)\cos(100x)$
@@
In this section, we demonstrate the performance of our methods compared to each baseline on 1D problems including the 1D Poisson, 1D steady-state Allen-Cahn, and 1D One-Way Wave equation. We first tested a set of 1D Poisson problems (see equation 12) with the following ground-truth solutions,
@@
The boundary conditions are derived from the true solution. We visualize the relative $L_{2}$ errors for each baseline method on these problems in Figures 4, 5, and 6 respectively. For Fourier PINNs, we set $K=150,200,250$ and $300$ for problems equation 29 and equation 30, then simply set $K=200$ for the remaining tests. Then, we examined the 1D steady-state Allen-Cahn equation of the form 14 with the same true manufactured solutions as for the 1D Poisson problem and report the results in Figures 7, 8, and 9. The final 1D problem we tested is the 1D One-Way Wave equation of the form,
@@
The results are reported in Figure 10.
Figure 7: 1D Allen-Cahn equation with true solution $u(x)=\sin(100x)$
Figure 8: 1D Allen-Cahn with true multi-scale solution $u(x)=\sin(x)+0.1\sin(20x)+0.05\cos(100x)$
Figure 9: 1D Allen-Cahn equation with true solution $u(x)=\sin(6x)\cos(100x)$
Figure 10: 1D One-Way Wave equation with true solution $u(x,t)=\sin(\pi x)\cos(10\pi t)+\sin(2\pi x)\cos(20\pi t)$
Figure 11: 2D Poisson with $u(x,y)=\sin(100x)\sin(100y)$
Figure 12: 2D Poisson with $u(x,y)=\sin(6x)\cos(20x)+\sin(6y)\cos(20y)$
Figure 13: 2D Allen-Cahn with $u(x,y)=(\sin(x)+0.1\sin(20x)+\cos(100x))\cdot(\sin(y)+0.1\sin(20y)+\cos(100y))$
6.4 2D Numerical Experiments
In this section, we demonstrate the performance of our methods compared to each baseline on 2D problems, including the 2D Poisson and 2D steady-state Allen-Cahn equations. We first tested on the 2D Poisson problem given by,
$\Delta u$ $=f(x,y),\;\;x,y\in[0,2\pi],$ (34)
with the following ground-truth manfactured solutions,
$u(x,y)$ $=\sin(100x)\sin(100y),$ (35)
$u(x,y)$ $=\sin(6x)\cos(20x)+\sin(6y)\cos(20y).$ (36)
We visualize the relative $L_{2}$ errors for each baseline method on these examples in Figures 11 and 12 respectively. We next tested on the 2D steady-state Allen-Cahn problem which is given by.
$u_{xx}+u_{yy}+u(u^{2}-1)$ $=f(x,y),\;\;x,y\in[0,2\pi],$ (38)
with the ground-truth solution,
$u(x,y)$ $=\left(\sin(x)+0.1\sin(20x)+\cos(100x)\right)$
$\;\;\;\cdot\left(\sin(y)+0.1\sin(20y)+\cos(100y)\right),$ (39)
and we report the results in Figure 13. Note that in these tests, the RFF-PINN method failed and resulted in large relative errors for all the single-scale settings, therefore we did not report these results.

### v2

@@ 6 Experiments
This section comprehensively evaluates our methods across a diverse set of partial differential equation (PDE) problems. The test cases include the 1D and 2D Poisson equations, which serve as canonical examples of linear elliptic PDEs, and the 1D and 2D steady-state Allen-Cahn equations, representing nonlinear elliptic PDEs with a nonlinear reaction term. To further challenge the models, we analyze the 1D one-way Wave equation, which incorporates a first-order time derivative, and the 1D non-steady-state Allen-Cahn equation, a stiff PDE with non-harmonic characteristics. We first describe the baseline methods for comparing Fourier PINNs and then outline the experimental setup. Finally, we present and discuss the results, highlighting the relative performance of each method on the benchmark problems.
@@ 6.1 Baselines
We benchmarked against the following state-of-the-art PINN models for solving high-frequency and multi-scale solutions:
@@
    (RFF-PINN) Random Fourier Feature PINNs (Wang et al., 2021c) with dynamically re-weighted loss terms derived from the NTK eigenvalues as in Wang et al. (2022),
@@
    (W-PINN) Weighted PINNs that down-weight the residual loss term to reduce its dominance and to fit the boundary loss Wang et al. (2022) better;
@@
    (Spectral) A single Fourier layer comprised of Fourier bases and trainable coefficients. Basis coefficients are optimized through the same gradient descent routines as all other baselines [^3] Throughout the experiments, we used the same set of Fourier bases for the spectral method and Fourier PINNs..
We do not test against the conceptually similar Fourier Neural Operators (FNOs) (Li et al., 2021), as they are designed for inverse problems and rely on a data loss term. In contrast, our method focuses on solving forward problems and does not require training data within the domain. Additionally, while the methods listed above represent widely-used PINN architectures, we acknowledge the existence of alternative neural network approaches based on random features and random projections, such as Random Vector Functional Link networks (RVFLNs) (Zhang & Suganthan, 2016), Extreme Learning Machines (ELMs) (Ding et al., 2013; Fabiani et al., 2021; Calabrò et al., 2021), Random Projection Neural Networks (RPNNs) (Rahimi & Recht, 2008), Reservoir Computing (Pathak et al., 2018), and recent advancements like RandONet (Fabiani et al., 2025). While methods such as RandONets and RVFLNs demonstrate significant computational efficiency and excel in inverse problem-solving tasks, their ability to address spectral bias in forward high-frequency problems remains underexplored.
By contrast, the proposed Fourier PINN architecture directly targets spectral bias—a key limitation of existing PINN approaches in learning high-frequency and multi-scale solutions. Unlike conventional architectures, the Fourier PINN framework is both modular and adaptable, with the integration of Fourier bases capable of enhancing or complementing any of these alternative methods. Given its flexibility and minimal downsides, exploring how Fourier PINNs could be combined with these approaches to improve scalability and accuracy remains an exciting direction for future work.
Table 2: The number and corresponding scales of Gaussian variances utilized in the RFF-PINN. Scales were selected randomly from a set range of $[1,200]$. Preference was given to the subset $[1,20,50,100]$ to ensure a balanced distribution across the available range.
@@ 6.2 Hyperparameters and experimental details
For each experiment, we trained the models in two stages: first using the Adam optimizer (Kingma & Ba, 2015), followed by L-BFGS optimization (Liu & Nocedal, 1989) until convergence, with the tolerance set to $10^{-9}$. Specific hyperparameters and training details are provided in the following tables: Table 5 for the 1D Poisson and 1D Allen-Cahn experiments, Table 6 for the 1D Wave equation experiments, Table 7 for the 2D Poisson and 2D (steady-state) Allen-Cahn experiments, and Table 8 for the 1D (non-steady-state) Allen-Cahn experiment. All experiments use 32-bit floating-point precision (float32) to ensure computational efficiency and consistency across models and were conducted on a GeForce RTX 3090 GPU with CUDA version 12.3, running on Ubuntu 20.04.6 LTS. In Fourier PINNs, we determined the range of frequency candidates for the Fourier bases by setting a maximum frequency $K$ and including all equally spaced frequencies in the set $\{1,2,...,K\}$. For the alternating optimization algorithm (Algorithm 1), we set the basis truncation threshold $\delta=10^{-4}$ and the regularization strength parameter $\alpha=10^{-4}$. Optimization used the same number of Adam iterations across experiments, with pauses at every 1000th iteration, to solve the LS problem for five iterations. RFF-PINNs required specifying the number and scales of Gaussian variances used to construct the random features. To evaluate the effects of these hyperparameters, we tested 20 different settings, varying the number of scales (one, two, three, and five) and using variances suggested in (Wang et al., 2021c) along with randomly sampled values. Table 2 provides the exact sets of scales. For W-PINN, we varied the weight of the residual loss from $\{10^{-1},10^{-3},10^{-4}\}$ and reported the configuration achieving the best results. For A-PINN, we introduced a learned parameter for the activation function in each layer and updated these parameters jointly. We ran every method for five random trials and reported the five number summaries obtained on seperate testing datasets for each experiment in boxplot form. All code is implemented using the PyTorch C++ library Paszke et al. (2019) [^4] Code is available at https://github.com/VarShankar/KernelPack/tree/sciml.
@@
This section evaluates Fourier PINNs on three progressively challenging 1D Poisson and Allen-Cahn benchmarks, demonstrating their superior performance and robustness over the baseline methods. These benchmarks systematically assess the models’ ability to handle high-frequency, multi-scale, and hybrid-frequency solutions for linear and non-linear elliptic PDEs. The analysis includes relative $\ell_{2}$ error comparisons, convergence dynamics, and frequency spectrum decomposition to highlight the strengths of Fourier PINNs. Table 5 in the Appendix lists each example’s hyperparameters and experimental details.
6.3.1 1D Poisson Problem
The 1D Poisson problem evaluates the ability of Fourier PINNs to capture high-frequency signals within a linear context. We test the following manufactured solutions with increasing frequency complexity:
@@
The boundary conditions are derived directly from the true solutions. These tests progressively introduce multi-scale and hybrid frequency components used to study the robustness and scalability of each model.
Figure 7: Results for the single-scale true solution $u(x)=\sin(100x)$. (a) Boxplot showing the relative $\ell_{2}$ errors for Fourier PINNs and the top-performing baseline methods. (b) Convergence plots comparing the relative $\ell_{2}$ error across iterations (top) and training time (bottom) for selected models. Vertical lines indicate when each model reached an error of $10^{-2}$. (c) Frequency spectrum of the solution $u(x)$ compared to PINN and Fourier PINN approximations. (d) Decomposition of the frequency components in Fourier PINNs between the NN output and the Fourier layer output, highlighting differences in spectral representation.
Single-Scale Solution Results. Figure 7(a) shows the relative $\ell_{2}$ error distributions for Fourier PINNs and top baseline models, while Figure 15(a) in the Appendix includes results for all models. Fourier PINNs consistently achieve significantly lower errors than baselines, including RFF-PINNs and spectral methods. Notably, RFF-PINNs demonstrate sensitivity to the chosen scale configurations, which impacts their overall performance. Figure 7(b) illustrates the convergence behavior across iterations and training time to analyze training dynamics. Fourier PINNs exhibit rapid and stable convergence due to the alternating optimization and basis truncation routine. In contrast, standard PINNs plateau at higher errors, struggling to approximate the high-frequency features of the solution accurately.
As shown in Figure 7(c), the frequency domain analysis compares the Discrete Fourier Transform (DFT) of the solutions from standard PINNs and Fourier PINNs against the ground truth. Fourier PINNs generate frequency spectra almost identical to the ground truth, effectively capturing the high-frequency features. Conversely, standard PINNs fail to resolve the high-frequency components, underscoring their limitations for high-frequency problems. To further dissect Fourier PINN solutions, we decomposed the predicted solution, $u_{F}=u_{N}+u_{B}$, into its neural network (NN) and Fourier layer components. Figure 7(d) presents the frequency spectra of each component. The absence of green bars (representing the NN component) indicates no contribution from the NN basis. This result aligns with the basis truncation outcome, where all NN bases were truncated during training, leaving only a single Fourier basis. This observation confirms that the optimization algorithm accurately captures the frequency components of the true solution.
Figure 8: (1D Poisson) Results for the multi-scale true solution $u(x)=\sin(x)+0.1\sin(20x)+0.05\cos(100x)$. (a) Boxplot showing the relative $\ell_{2}$ errors for Fourier PINNs and the top-performing baseline methods. (b) Convergence plots comparing the relative $\ell_{2}$ error across iterations (top) and training time (bottom) for selected models. Vertical lines indicate when each model reached an error of $10^{-3}$. (c) Frequency spectrum of the solution $u(x)$ compared to PINN and Fourier PINN approximations. (d) Decomposition of the frequency components in Fourier PINNs between the NN output and the Fourier layer output, highlighting differences in spectral representation.
Multi-Scale Solution Results. The multi-scale problem introduces a combination of low- and high-frequency components, presenting a greater challenge for baseline methods. Figure 8(a) shows the relative $\ell_{2}$ error distributions for Fourier PINNs and top baseline models, while Figure 15(b) in the Appendix includes results for all models. Fourier PINNs outperform most baseline methods; however, a subset of RFF-PINN configurations achieves slightly better results. These cases highlight the sensitivity of RFF-PINNs to scale selection, whereas Fourier PINNs demonstrate robust performance across all tested basis configurations. RFF-PINNs exhibit more significant variability across configurations, with inconsistent performance due to their reliance on precise scale tuning.
The frequency spectrum comparison in Figure 8(c) illustrates that Fourier PINNs accurately reconstruct both low- and high-frequency components, highlighting their adaptability to multi-scale structures. In contrast, the spectrum of the standard PINN output shows that it accurately captures low-frequency components but struggles to resolve the solution’s high-frequency features. Figure 8(d) further analyzes the frequency decomposition of the Fourier PINN solution. The Fourier layer captures all high-frequency components of the target solution (e.g., $\sin(20x)$ and $\cos(100x)$ in Equation 35), while the neural network (NN) output contributes to modeling the low-frequency component ($\sin(x)$). This result aligns with the basis truncation outcomes: three Fourier bases were retained, and approximately 10% of the NN basis functions were truncated during training. These findings demonstrate that the hybrid architecture of Fourier PINNs, combined with the alternating optimization and truncation algorithm, effectively balances the modeling of both smooth and oscillatory components in multi-scale solutions.
Figure 9: (1D Poisson) Results for the single-scale true solution $u(x)=\sin(6x)\cos(100x)$. (a) Boxplot showing the relative $\ell_{2}$ errors for Fourier PINNs and the top-performing baseline methods. (b) Convergence plots comparing the relative $\ell_{2}$ error across iterations (top) and training time (bottom) for selected models.
Hybrid Solution Results. This hybrid-frequency example introduces closely spaced high-frequency components, which increases the complexity of disentangling and modeling each frequency. Figure 9(a) illustrates the relative $\ell_{2}$ error distributions for Fourier PINNs and top baseline models, while Figure 15(c) in the Appendix presents results for all models. To specifically evaluate the impact of the alternating optimization algorithm on Fourier PINNs, we tested multiple basis configurations using the alternating optimization method described in Algorithm 1, as well as traditional optimization routines (i.e., without the least-squares (LS) basis coefficient optimization).
The results demonstrate that Fourier PINNs, when combined with the alternating optimization strategy, consistently achieve the lowest $\ell_{2}$ errors across all tested configurations, highlighting their ability to effectively model solutions with coupled and closely spaced frequency components. In contrast, Fourier PINNs without alternating optimization fail to accurately capture the true solution, emphasizing the critical role of the LS basis coefficient updates in resolving hybrid-frequency structures. The convergence behavior shown in Figure 9(b) further underscores the efficiency of Fourier PINNs in learning hybrid solutions. Their rapid convergence and stable performance across iterations showcase their robustness and adaptability to challenging problem setups with complex frequency interactions. Additionally, there is little variation in training time with increasing Fourier basis sizes.
6.3.2 1D Steady-State Allen-Cahn Problem
Next, we examine the 1D steady-state Allen-Cahn equation, which is a nonlinear reaction-diffusion system. This problem serves as a more complex PDE benchmark for evaluating Fourier PINNs and baseline models. To ensure consistency, we test the same manufactured solutions used in the 1D Poisson case, defined by Equation 34, Equation 35, and Equation 36.
Results. Figure 16 presents the relative $\ell_{2}$ error distributions for Fourier PINNs and the top-performing baseline models across all manufactured solutions, while detailed results for all models can be found in Figure 16 in the Appendix. Across all configurations, Fourier PINNs consistently achieve lower relative $\ell_{2}$ errors. Additionally, their faster convergence, as highlighted in Figure 11, underscores the efficiency of the Fourier PINN framework for solving the 1D steady-state Allen-Cahn problem.
Figure 10: (1D Allen-Cahn) Boxplot of relative $\ell_{2}$ errors for Fourier PINNs and top-performing baseline methods across the manufactured solutions: left, $u(x)=\sin(100x)$; middle, $u(x)=\sin(x)+0.1\sin(20x)+0.05\cos(100x)$; and right, $u(x)=\sin(6x)\cos(100x)$.
Figure 11: Convergence behavior of select models on the 1D steady-state Allen-Cahn problems, showing the mean relative $\ell_{2}$ error as a function of iterations (left) and training time (right). Results are presented for $u(x)=\sin(100x)$ (left two plots) and $u(x)=\sin(x)+0.1\sin(20x)+0.05\cos(100x)$ (right two plots).
6.4 2D Numerical Experiments
We evaluate Fourier PINNs on two 2D PDEs: the 2D Poisson problem and the steady-state Allen-Cahn equation. These examples test the model’s scalability, ability to capture high-frequency components, and approximation of multi-scale solutions in higher dimensions. Table 7 in the Appendix lists each example’s hyperparameters and experimental details.
6.4.1 2D Poisson Problem
The 2D Poisson problem is given by:
$\Delta u$ $=f(x,y),\;\;x,y\in[0,2\pi],$ (37)
with the following ground-truth solutions:
$u(x,y)$ $=\sin(100x)\sin(100y),$ (38)
$u(x,y)$ $=\sin(6x)\cos(20x)+\sin(6y)\cos(20y).$ (39)
Figure 12: Convergence behavior of select models on the 2D Poisson problems, showing the mean relative $\ell_{2}$ error as a function of iterations (left) and training time (right). Results are presented for Equation 38 (left two plots) and Equation 39 (right two plots).
Results. For the high-frequency solution (Equation 38), Fourier PINNs achieve an average relative $\ell_{2}$ error of 0.0384, outperforming all baseline methods, with the spectral method being the closest competitor at 0.0423. The Appendix provides detailed results for all models in Figure 17. Notably, other methods fail to capture the high-frequency features, with errors exceeding 1. For the multi-scale solution (Equation 39), Fourier PINNs demonstrate a significant advantage, achieving an average relative $\ell_{2}$ error of 0.00085, far surpassing the performance of all other baselines. Figure 12 highlights the superior convergence behavior of Fourier PINNs in both scenarios, underscoring their ability to handle challenging high-frequency and multi-scale problems effectively.
6.4.2 2D Steady-State Allen-Cahn Equation
The Allen-Cahn problem is defined as:
$u_{xx}+u_{yy}+u(u^{2}-1)$ $=f(x,y),\;\;x,y\in[0,2\pi],$ (40)
with the ground-truth solution:
$u(x,y)$ $=\left(\sin(x)+0.1\sin(20x)+\cos(100x)\right)\cdot\left(\sin(y)+0.1\sin(20y)+\cos(100y)\right).$ (41)
Results: The Allen-Cahn example highlights Fourier PINNs’ challenges with nonlinear PDEs. The model achieves a relative $\ell_{2}$ error of 0.9, outperforming all other methods ($>1$). In the Appendix, Figure 18 provides detailed results for all models. This suggests the need for further optimization for highly nonlinear problems.
Figure 13: (1D One-Way Wave) Boxplot of relative $\ell_{2}$ errors for Fourier PINNs and top-performing baseline methods.
6.5 1D One-Way Wave Problem
We next tested the 1D One-Way Wave equation of the form,
@@
This problem introduces a time-dependent component, testing the ability of models to handle first-time derivatives within a high-frequency context. We conducted an ablation study to assess the roles of basis pruning and regularization by evaluating Fourier PINNs without these mechanisms. This study highlights the distinct contributions of basis pruning and regularization to model accuracy and computational efficiency. Table 6 in the Appendix lists all hyperparameters and experimental details.
Figure 13 shows the relative $\ell_{2}$ error distributions for Fourier PINNs and top-performing baseline models across all manufactured solutions. Detailed results for all models are provided in Figure 19 in the Appendix. As expected, Fourier PINNs with both basis regularization and pruning (denoted as "reg + prune") achieve the best results, while the configuration using neither (denoted as "no reg + no prune") yields the worst performance.
Fourier-PINN Without Basis Pruning. In configurations without basis pruning, all frequency candidates up to the maximum frequency $K$ were allowed to contribute to the solution. Two variations were tested: with basis regularization ("reg + no prune") and without regularization ("no reg + no prune"). Both configurations resulted in worse relative $\ell_{2}$ errors than the pruned counterparts, although the regularized version performed slightly better than the unregularized one.
Fourier-PINN Without Regularization. To further evaluate the impact of regularization, we tested configurations without the regularization term. Two setups were compared: with basis pruning ("no reg + prune") and without pruning ("no reg + no prune"). The configuration without regularization but with basis pruning ("no reg + prune") achieved better relative $\ell_{2}$ errors than the setup using regularization but no basis pruning ("reg + no prune"). This result suggests that basis pruning has a slightly more significant influence on solution quality than regularization.
6.6 1D Non-Steady-State Allen-Cahn Problem
The time-dependent Allen-Cahn equation is a well-established benchmark for testing the limitations of PINN models, particularly due to its stiffness, sharp transitions, and nonlinear dynamics (Wight & Zhao, 2020; Wang et al., 2024b). For simplicity, we focus on a one-dimensional case with periodic boundary conditions over $t\in[0,1]$ and $x\in[-1,1]$:
$u_{t}-0.0001u_{xx}+5u^{3}-5u=0,$
$u(0,x)=x^{2}\cos(\pi x),$
$u(t,-1)=u(t,1),\;u_{x}(t,-1)=u_{x}(t,1).$
This problem is particularly challenging because the Allen-Cahn PDE introduces complex, multi-scale dynamics driven by a combination of stiff reaction terms and spatial diffusion. Unlike problems with harmonic solutions, the nonlinear terms generate sharp transitions and localized structures in both space and time, testing the model’s adaptability to varying frequency components and its robustness to sharp gradients. Table 8 in the Appendix outlines the hyperparameter and experimental settings used in the example. In this experiment, we randomly sampled $8192$ collocation points within the domain at each iteration. Figure 14 compares the convergence behavior of Fourier PINNs with baseline models (standard PINN and RFF-PINN).
Figure 14: Convergence behavior of different models on the 1D non-steady-state Allen-Cahn equation. Fourier PINNs achieve superior accuracy and convergence, effectively capturing sharp transitions and localized structures, while baseline methods falter.
Results. As shown in Figure 14, the Fourier PINN significantly outperforms both the standard PINN and RFF-PINN. The relative $\ell_{2}$ error decreases steadily for Fourier PINNs, reaching values below $10^{-3}$ after sufficient training iterations, while the baseline models plateau at higher error levels. Specifically, the standard PINN struggles with stiffness, failing to reduce the error below $10^{-1}$ even after extended training. The RFF-PINN shows some improvement over the standard PINN but is highly sensitive to scale settings, resulting in slower convergence and larger errors compared to Fourier PINNs. Fourier PINNs leverage the augmented Fourier layer structure to efficiently represent both smooth and sharp features, as evidenced by their rapid convergence and low final error. This experiment underscores the potential of Fourier PINNs to handle challenging nonlinear PDEs with sharp transitions, where conventional PINN architectures struggle.